OpenStudy (anonymous):

Which of the following situations could be represented with the given system of equations? 0.1x+0.6y=0.4(5) x+y=5 A. Eunice needs to combine solutions of 0.1M HCl and 0.4M HCl to produce 5L of 0.6M hydrochloric acid (HCl). B. Eunice needs to combine solutions of 0.1M HCl and 0.6M HCl to produce 5L of 0.4M hydrochloric acid (HCl). C. Eunice needs to combine solutions of 0.4M HCl and 0.6M HCl to produce 5L of 0.1M hydrochloric acid (HCl). D. Eunice needs to combine solutions of 0.1M HCl and 0.6M HCl to produce 4L of 0.5M hydrochloric acid (HCl).

OpenStudy (anonymous):

@joquez14 help

OpenStudy (anonymous):

@Megadevon

OpenStudy (anonymous):

@xxkuroshinigamixx

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@watercolour_

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@SarahEZZMcK

OpenStudy (anonymous):

@Karmelo1

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@Awesome781

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@anonymous97

OpenStudy (anonymous):

@WeMaejor900

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (anonymous):

@tgawade

OpenStudy (anonymous):

@adrynicoleb

OpenStudy (anonymous):

@Ness9630

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@kelliee

OpenStudy (e.mccormick):

Well, which do you think and why?

OpenStudy (anonymous):

Honestly I don't get this question I need all the help. Which do you think?

OpenStudy (e.mccormick):

I know the answer. With a system like this, you have two concepts that are working together to make the answers. The first concept is that two volumes of different concentrations are added to make a new concentration. The other is that two volumes are added to make a total volume.

OpenStudy (anonymous):

uhm

OpenStudy (anonymous):

I think it's B.. not sure

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

nevermind, thanks for your help.

OpenStudy (e.mccormick):

OK, lets look. 0.1x+0.6y=0.4(5) x+y=5 B. Eunice needs to combine solutions of 0.1M HCl and 0.6M HCl to produce 5L of 0.4M hydrochloric acid (HCl). Yes. If x is the 0.1M HCl and y is the 0.6M HCl then: 0.1x+0.6y=0.4 \(\leftarrow\) the 0.1 and 0.6 mixed make 0.4 x+y=5 \(\leftarrow\) the total volumes should add to 5.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I kinda understand now

OpenStudy (anonymous):

still need help sorry i was away